# BRIGHT: A globally distributed multimodal building damage assessment dataset with very-high-resolution for all-weather disaster response

### v2

@@ Appendix G Introduction of UMCD methods and a more practical evaluation protocol
Hongruixuan Chen: conceptualization (lead), data curation (lead), funding acquisition (equal), methodology (lead), project administration (lead), investigation (lead), software (lead), visualization (lead), writing – original draft preparation (lead), writing – review and editing (lead). Jian Song: conceptualization (support), data curation (equal), funding acquisition (support), methodology (support), investigation (support), software (support), writing – original draft preparation (equal), visualization (support). Olivier Dietrich: conceptualization (support), data curation (equal), software (support), visualization (equal), writing – review and editing (support). Clifford Broni-Bediako: conceptualization (support), data curation (support), writing – review and editing (equal). Weihao Xuan: data curation (support), project administration (support), writing – review and editing (equal). Junjue Wang: data curation (support), writing – review and editing (support). Xinlei Shao: funding acquisition (support), writing – review and editing (support). Yimin Wei: writing – review and editing (support). Junshi Xia: conceptualization (support), data curation (support), supervision (support), writing – review and editing (support). Cuiling Lan999Cuiling Lan did not participate in any activities related to the acquisition, processing, utilization, or distribution of the datasets. Her access to the data was solely limited to the information presented in the current paper, similar to that of the readers. : conceptualization (support), supervision (support), writing – linguistic refinement (equal). Konrad Schindler: conceptualization (equal), resources (support), supervision (equal), writing – review and editing (equal). Naoto Yokoya: conceptualization (equal), funding acquisition (lead), resources (lead), supervision (lead), writing – review and editing (equal).
@@ Acknowledgements.
This work was supported in part by the JSPS, KAKENHI under Grant Number 24KJ0652 and 22H03609, the Council for Science, Technology and Innovation (CSTI), the Cross-ministerial Strategic Innovation Promotion Program (SIP), Development of a Resilient Smart Network System against Natural Disasters (Funding agency: NIED), JST, FOREST under Grant Number JPMJFR206S, Microsoft Research Asia, Next Generation AI Research Center of The University of Tokyo, and Young Researchers Exchange Programme between Japan and Switzerland under the Japanese-Swiss Science and Technology Programme. The authors would also like to give special thanks to Sarah Preston of Capella Space, Capella Space’s Open Data Gallery, Maxar Open Data Program, and Umbra Space’s Open Data Program for providing the valuable data.References

### v3

@@ Appendix H Details of unsupervised multimodal image matching methods
Hongruixuan Chen: conceptualization (lead), data curation (lead), funding acquisition (equal), methodology (lead), project administration (lead), investigation (lead), software (lead), visualization (lead), writing – original draft preparation (lead), writing – review and editing (equal). Jian Song: conceptualization (support), data curation (equal), funding acquisition (support), methodology (support), investigation (support), software (support), writing – original draft preparation (equal), visualization (support). Olivier Dietrich: conceptualization (support), data curation (equal), software (support), visualization (equal), writing – review and editing (support). Clifford Broni-Bediako: conceptualization (support), data curation (support), writing – review and editing (lead). Weihao Xuan: data curation (support), project administration (support), writing – review and editing (equal). Junjue Wang: data curation (support), writing – review and editing (support). Xinlei Shao: funding acquisition (support), writing – review and editing (support). Yimin Wei: writing – review and editing (support). Junshi Xia: conceptualization (support), data curation (support), supervision (support), writing – review and editing (support). Cuiling Lan121212Cuiling Lan did not participate in any activities related to the acquisition, processing, utilization, or distribution of the datasets. Her access to the data was solely limited to the information presented in the current paper, similar to that of the readers. : conceptualization (support), supervision (support), writing – linguistic refinement (equal). Konrad Schindler: conceptualization (equal), resources (support), supervision (equal), writing – review and editing (equal). Naoto Yokoya: conceptualization (equal), funding acquisition (lead), resources (lead), supervision (lead), writing – review and editing (equal).
@@ Acknowledgements.
This work was supported in part by the JSPS, KAKENHI under Grant Number 24KJ0652 and 22H03609, the Council for Science, Technology and Innovation (CSTI), the Cross-ministerial Strategic Innovation Promotion Program (SIP), Development of a Resilient Smart Network System against Natural Disasters (Funding agency: NIED), JST, FOREST under Grant Number JPMJFR206S, Microsoft Research Asia, Next Generation AI Research Center of The University of Tokyo, and Young Researchers Exchange Programme between Japan and Switzerland under the Japanese-Swiss Science and Technology Programme. The authors would also like to give special thanks to Sarah Preston of Capella Space, Capella Space’s Open Data Gallery, Maxar Open Data Program, and Umbra’s Open Data Program for providing the valuable data.References
@@
- Andrienko et al. (2025) Andrienko, D., Goriunov, D., Grudova, V., Markuts, J., Marshalok, T., Neyter, R., Piddubnyi, I., Studennikova, I., and Topolskov, D.: Report on Damages to Infrastructure from the Destruction Caused by Russia’s Military Aggression Against Ukraine as of November 2024, Tech. rep., Kyiv School of Economics (KSE) Institute, 2025.
@@
- Dietrich et al. (2025) Dietrich, O., Peters, T., Garnot, V. S. F., Sticher, V., Whelan, T. T.-T., Schindler, K., and Wegner, J. D.: An Open-Source Tool for Mapping War Destruction at Scale in Ukraine Using Sentinel-1 Time Series, Communications Earth & Environment, 6, 215, 2025.
- Fakih and Majzoub (2021) Fakih, L. and Majzoub, A.: "They Killed Us from the Inside": An Investigation Into the August 4 Beirut Blast, Human Rights Watch, 2021.
@@
- Fire and Disaster Management Agency (2025) Fire and Disaster Management Agency: Damage Caused by the 2024 Noto Peninsula Earthquake and Response Status of Fire Departments, etc. (117th Report), Situation Report 117, Fire and Disaster Management Agency, Tokyo, Japan, 2025.
- Forrister et al. (2024) Forrister, A., Kuligowski, E. D., Sun, Y., Yan, X., Lovreglio, R., Cova, T. J., and Zhao, X.: Analyzing Risk Perception, Evacuation Decision and Delay Time: A Case Study of the 2021 Marshall Fire in Colorado, Travel Behaviour and Society, 35, 100 729, 2024.
@@
- GFDRR and World Bank (2023) GFDRR and World Bank: Extremely Severe Cyclonic Storm Mocha, May 2023, Myanmar: Global Rapid Post-Disaster Damage Estimation (GRADE) Report, Tech. rep., Global Facility for Disaster Reduction and Recovery (GFDRR) and World Bank, 2023.
@@
- Government of Türkiye (2023) Government of Türkiye: Türkiye Earthquakes Recovery and Reconstruction Assessment, Tech. rep., Government of Türkiye, https://reliefweb.int/report/turkiye/turkiye-earthquakes-recovery-and-reconstruction-assessment, 2023.
@@
- Hedayati et al. (2024) Hedayati, F., Monroy, X., Sluder, E., Fallahian, H., and Shabanian, M.: The 2023 Lahaina Conflagration, https://ibhs.org/wp-content/uploads/FINAL-Lahaina-Conflagration.pdf, 2024.
@@
- IFRC (2021) IFRC: Emergency Appeal: Democratic Republic of Congo and Rwanda, Report, International Federation of Red Cross and Red Crescent Societies (IFRC), https://reliefweb.int/report/democratic-republic-congo/emergency-appeal-democratic-republic-congo-and-rwanda-mount, 2021.
- International Federation of Red Cross and Red Crescent Societies (2024) International Federation of Red Cross and Red Crescent Societies: Morocco Earthquake 2023: Operation Update #4 (12 months) (MDRMA010), Situation Report MDRMA010OU5, International Federation of Red Cross and Red Crescent Societies, Geneva, https://reliefweb.int/report/morocco/morocco-earthquake-2023-operation-update-4-12-months-mdrma010, published 16 November 2024; accessed 18 April 2025, 2024.
- Japanese Red Cross Society (2024) Japanese Red Cross Society: Operation Update No. 30: 2024 Noto Peninsula Earthquake—The Japanese Red Cross Society’s Response, https://www.jrc.or.jp/english/relief/2024NotoPeninsulaEarthquake.html, web page, 2024.
- Jones et al. (2024) Jones, M. W., Kelley, D. I., Burton, C. A., Di Giuseppe, F., Barbosa, M. L. F., Brambleby, E., Hartley, A. J., Lombardi, A., Mataveli, G., McNorton, J. R., Spuler, F. R., Wessel, J. B., Abatzoglou, J. T., Anderson, L. O., Andela, N., Archibald, S., Armenteras, D., Burke, E., Carmenta, R., Chuvieco, E., Clarke, H., Doerr, S. H., Fernandes, P. M., Giglio, L., Hamilton, D. S., Hantson, S., Harris, S., Jain, P., Kolden, C. A., Kurvits, T., Lampe, S., Meier, S., New, S., Parrington, M., Perron, M. M. G., Qu, Y., Ribeiro, N. S., Saharjo, B. H., San-Miguel-Ayanz, J., Shuman, J. K., Tanpipat, V., van der Werf, G. R., Veraverbeke, S., and Xanthopoulos, G.: State of Wildfires 2023–2024, Earth System Science Data, 16, 3601–3685, 2024.
- Kaneko (2024) Kaneko, Y.: Prioritizing Recovery: Expected Roles of Public Support during the Noto Peninsula Earthquake, Research Report of the Research Center for Urban Safety and Security, Kobe University, 28, 31–39, http://www.rcuss.kobe-u.ac.jp/publication/Year2024/pdfEach28/28_05.pdf, article in English; RCUSS Research Report No. 28, 2024.
@@
- National Geophysical Data Center / World Data Service (2025) National Geophysical Data Center / World Data Service: NCEI/WDS Global Significant Earthquake Database, 10.7289/V5TD9V7K, accessed: 2025-04-18, 2025.
- NCEI (2025) NCEI: U.S. Billion-Dollar Weather and Climate Disasters, Tech. rep., NOAA National Centers for Environmental Information (NCEI), https://www.ncei.noaa.gov/access/billions/, 2025.
@@
- Normand and Heggy (2024) Normand, J. C. and Heggy, E.: Assessing flash flood erosion following storm Daniel in Libya, Nature Communications, 15, 6493, 2024.
- OCHA (2021a) OCHA: Equatorial Guinea: Bata Explosions - Flash Update No. 3 (As of 11 March 2021), Report, United Nations Office for the Coordination of Humanitarian Affairs (OCHA), https://reliefweb.int/report/equatorial-guinea/equatorial-guinea-bata-explosions-flash-update-no3-11-march-2021, 2021a.
- OCHA (2021b) OCHA: Haiti: Earthquake Situation Report No. 4, Report, United Nations Office for the Coordination of Humanitarian Affairs (OCHA), https://reliefweb.int/report/haiti/haiti-earthquake-situation-report-no-4-7-september-2021, 2021b.
- OCHA (2023) OCHA: Morocco Earthquake: Situation Report #5, Tech. rep., United Nations Office for the Coordination of Humanitarian Affairs (OCHA), https://reliefweb.int/report/morocco/morocco-earthquake-situation-report-5-november-14-2023, 2023.
@@
- Reinhart and Reinhart (2024) Reinhart, B. J. and Reinhart, A.: Hurricane Otis (EP182023): Tropical Cyclone Report, Tropical Cyclone Report EP182023, National Hurricane Center, National Oceanic and Atmospheric Administration, Miami, FL, https://www.nhc.noaa.gov/data/tcr/EP182023_Otis.pdf, issued 7 March 2024; PDF archived and retrieved the same day, 2024.
@@
- STL (2024) STL: Turkey Earthquake: Emergency Situation Report (05.02.2024), Tech. rep., Support to Life (STL), https://reliefweb.int/report/turkiye/turkey-earthquake-emergency-situation-report-05022024, 2024.
@@
- Troll et al. (2024) Troll, V. R., Aulinas, M., Carracedo, J. C., Geiger, H., Perez-Torrado, F. J., Soler, V., Deegan, F. M., Bloszies, C., Weis, F., Albert, H., Gisbert, G., Day, J. M. D., Rodríguez-Gonzalez, A., Gazel, E., and Dayton, K.: The 2021 La Palma eruption: social dilemmas resulting from life close to an active volcano, Geology Today, 40, 96–111, 2024.
@@
- USAID Bureau for Humanitarian Assistance (2024) USAID Bureau for Humanitarian Assistance: Libya Assistance Overview, Fact sheet, U.S. Agency for International Development, Washington, DC, https://reliefweb.int/report/libya/libya-assistance-overview-april-2024, pDF fact sheet (posted 17 April 2024), 2024.
- Valsamos et al. (2021) Valsamos, G., Larcher, M., and Casadei, F.: Beirut explosion 2020: A case study for a large-scale urban blast simulation, Safety Science, 137, 105 190, 2021.
@@
- Xia et al. (2025) Xia, J., Chen, H., Broni-Bediako, C., Wei, Y., Song, J., and Yokoya, N.: OpenEarthMap-SAR: A Benchmark Synthetic Aperture Radar Dataset for Global High-Resolution Land Cover Mapping, arXiv preprint arXiv:2501.10891, 2025.